.Using 8’s complement arithmetic compute the value of 6278-6758 and write your final answer in decimal
1
Expert's answer
2020-06-08T16:43:31-0400

1) convert Decimal to Octal:

6278/8 = 784(Quotient) and 6(Remainder)

784/8 = 98(Quotient) and 0(Remainder)

98/8 = 12(Quotient) and 2(Remainder)

12/8 = 1(Quotient) and 4(Remainder)

1/8 = 0(Quotient) and 1(Remainder)

X = 627810 = 142068

Y = 675810 = 151468


2) find 8's complement of the Y (8's complement of a number is 1 added to it's 7's complement number):

777778 - 151468 = 626318

626318 + 18 = 626328


3) now Add this 8's complement of Y to X:

142068 + 626328 = 770408


4) here there is no carry, then answer is the 8's complement of the sum obtained with a minus:

777778 - 770408 = 7378

7378 + 18 = 7408

result: -7408


5) convert Octal to Decimal:

-7408 = -(7*82+4*81+0*80) = -48010


6) check:

6278 - 6758 = -480


Answer: -480